A standard airbag has 130.0 g of sodium azide (NaN3).  How many grams of dangerous sodium metal would be produced? 


Show your stoichiometry work in the tables below, including all appropriate units, for full credit. 



1
Expert's answer
2021-04-21T03:54:23-0400

The decomposition of NaN3 is represented by the following equation:

2NaN3 --> 2Na + 3N2

Therefore,

"m(Na)=130.0g(NaN_3)\times\frac{1mol(NaN_3)}{65.01g(NaN_3)}\times\frac{2mol(Na)}{2mol(NaN_3)}\times\frac{22.99g(Na)}{1mol(Na)}=45.97g"


Answer: 45.97 g